OpenStudy (anonymous):

find the zeros of the functions below f(x)= x^2+6x+10 f(x)= 3x^2-3x+9 f(x)= x^2-2x-9 f(x)= x^2+10x=5

OpenStudy (anonymous):

repeated applications of the quadratic formula

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[ x^2+6x+10=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=6,c=10\]

OpenStudy (anonymous):

in this case \(b^2-4ac=36-4\times 10=-4\) so there is no real zero

OpenStudy (anonymous):

can u just give me the answers to the problems i posted it confusing me to figure it out please and thank you :)